OpenStudy (anonymous):

What is the surface area of a square based pyramid with a height of 7.2 feet, slant height of 8.4 feet and base length of 8.8 feet?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

do you know the formula ??

Nnesha (nnesha):

are you there ??

OpenStudy (anonymous):

yeah

Nnesha (nnesha):

okay do you know the formula ??

OpenStudy (anonymous):

yess

Nnesha (nnesha):

okay so what is formula ?

OpenStudy (anonymous):

V = 1/3Bh

Nnesha (nnesha):

that's for volume which is suppose to be \[\huge\rm v= \frac{ 1 }{ 3 }B^2h\]

Nnesha (nnesha):

formula for surface area is \[\huge\rm A= 2bs +b^2\] where s is slant height and b is base so replace variables by their values and solve for a

OpenStudy (anonymous):

i thought A = 1/2 lp + B

OpenStudy (anonymous):

..

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

wait a sec let me see few more websites

Nnesha (nnesha):

here is another one https://www.easycalculation.com/area/learn-square-pyramid.php

OpenStudy (anonymous):

oh

Nnesha (nnesha):

what did you get ?